solve using elimination method:
4x-2y=1
4x-12y=2
(answer is ordered pair integers or fractions)

Answer by checkley77(12569)

4x-2y=1
-(4x-12y=2) SUBTRACT)
------------------------
10Y=-1
Y=-1/10
Y=-.1 ANS.
4X-2*-.1=1
4X+.2=1
4X=1-.2
4X=.8
X=.8/4
X=.2 ANS.
PROOF:
4*.2-12*-.1=2
.8+1.2=2